OpenStudy (anonymous):

Two number cubes are rolled. What is the probability of rolling a sum of 9 or a sum that is even? I get confused because I don't know if each would be a separate probability, like independent or dependent and if I multiply them or if they are together, what would I do?

OpenStudy (anonymous):

can the sum be 9 and even at the same time?

OpenStudy (anonymous):

btw this has nothing whatsoever to do with "independent" or "dependent" because you are not multiplying because the sum cannot be both 9 and even, those events are called "disjoint" to find the probability of one or the other ADD

OpenStudy (anonymous):

Okay so, I would take the total number of outcomes, and then take the amount of outcomes that can be a sum of 9 or 7, which would be 2/6 or 1/3?

OpenStudy (anonymous):

for one dice

OpenStudy (anonymous):

it says "sum is 9 OR even" right?

OpenStudy (anonymous):

Wow I misread that, I thought it said 7, sorry. Long day

OpenStudy (anonymous):

what does it mean by sum though? You should know english is not my first language so I struggle a bit with word problems.

OpenStudy (anonymous):

yeah must have been sum is 7 or 9 would be \[\frac{6}{36}+\frac{4}{36}\]

OpenStudy (anonymous):

why the 36?

OpenStudy (anonymous):

it means "roll two dice, record the total" which is usually what you do when you roll two dice (although not always)

OpenStudy (anonymous):

Ah okay, like monopoly.

OpenStudy (anonymous):

the 36 because there are 36 elements in the sample space (set of all possible outcomes) when you roll two dice 6 outcomes possible for the first die, another 6 for the second \(6\times 6=36\) total

OpenStudy (anonymous):

yeah, or craps

OpenStudy (anonymous):

or even my favorite candy land

OpenStudy (anonymous):

Oh okay, so now I understand 6 / 26 and 4/ 36, so I would need to add them?

OpenStudy (anonymous):

IF the question was "what is the probability you roll at 7 or a 9" yes, but it is not

OpenStudy (anonymous):

Yeah yeah I understand that. So, since it isn't. It's asking for even numbers, which correct me if I'm wrong, this is not my first language but even means 2,4,6, . . etc right?

OpenStudy (anonymous):

right, it means the total (when you add) is 2,4,6,8,10 or 12

OpenStudy (anonymous):

don't think too hard about that probability, half the numbers are even

OpenStudy (anonymous):

I'm afraid my brain is shutting down or something. Let me just re read the problem. For some reason today has not been a very good math day. So, two dices, each one has 6, right? So, half the numbers are even, 2,4 and 6. So there are two, there should be 6/36 that can be even? (I'm sorry if I'm wrong)

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

i am going to attach a table hold on

OpenStudy (anonymous):

Okay, I'm just trying to visualize this

OpenStudy (anonymous):

The only possible thing that comes on my head is 6/12 for even numbers and 1/12 for hte nine. Oh wait let me look at that

OpenStudy (anonymous):

that table is supposed to represent the 36 possible outcomes when you roll two dice for example \((3,2)\) means the left die is a 6 and the top die is a 2, for a total of 5

OpenStudy (anonymous):

we can list the number of ways to roll a 9 on two dice, there are only 4 ways to do it \[(5,4),(4,5),(6,3),(3,6)\]

OpenStudy (anonymous):

since we believe the dice are fair, so all 36 outcomes are equally likely, we would say the probability you roll a total of 9 is \(\frac{4}{36}\)

OpenStudy (anonymous):

Okay. . . I sort of understand that

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

now you want to add that the probability you roll an even number (the total ls even)

OpenStudy (anonymous):

that probability you can compute by counting, but it is easier just to realize that half the numbers are even, so it must be \(\frac{1}{2}\)

OpenStudy (anonymous):

Okay, so, that would be me taking the 4/36 living it alone right? Or is the 1/2 a total separate number? It's just you said "you want to add" and I don't know whether you mean adding the numbers or as in "you want to count this factor in your thought process"

OpenStudy (anonymous):

leaving*

OpenStudy (anonymous):

mean you add the probabilities \[\frac{1}{2}+\frac{4}{36}\]

OpenStudy (anonymous):

okay good! Phew I thought I was lost again.

OpenStudy (anonymous):

This gets me 11/18

OpenStudy (anonymous):

yes looks good

OpenStudy (anonymous):

Phew, wow,I'm sorry for all the trouble. It has been a very long day. I am truly grateful for your help.

OpenStudy (whpalmer4):

@moonrose just a side note. in this problem, the two probabilities are disjoint — the number rolled is either odd (2,4,6,8) or a 9, but it is not both. Some problems like this you may have a situation where the two cases overlap. For example, what is the probability that the sum is even, or there is a 1 showing on one or both of the dice? There your work would be a bit more complicated because you can't just add the two probabilities together — doing that would count some rolls twice. 1 1 would be counted both as being an even sum AND as a roll where where a 1 is showing on one or both of the dice. Think about it later, when your brain has recovered :-)